# Performance Estimation in Binary Classification Using Calibrated Confidence

### v1

@@ 5 Discussion
It is not immediately obvious, how the methods for unsupervised accuracy estimation referenced in the Introduction could be adapted to estimate any other classification metric besides accuracy. As such, CBPE sets the baseline for unsupervised estimation of all binary classification metrics, which can be derived from the confusion matrix. Although CBPE comes with relatively strong theoretical guarantees, it also has its limitations. First and foremost, the theoretical guarantees hinge on the assumption of perfect calibration, which is not achievable in any real-life scenario. However, our experiments show, that in most cases, CBPE yields consistent estimates with relatively small estimation errors as the calibration errors remain small as well. This reduces the problem of performance estimation into a calibration error minimization problem.
Confidence-based performance estimators are known to struggle under concept shift, since even when $p_{s}(\boldsymbol{x},y)$ and $p_{t}(\boldsymbol{x})$ are fully known, the problem of estimating model performance in this setting is fundamentally unidentifiable [Garg et al., 2022] and no guarantees for estimation quality can be given. This is because concept shift breaks the stochastic dependency between the features and labels, rendering the confidence scores of a model obsolete in most cases. Also, since calibration error typically increases with covariate shift [Ovadia et al., 2019] and label shift [Popordanoska et al., 2023], methods for remaining calibrated under these shifts are needed. Unfortunately, no dedicated benchmarks for covariate shift and label shift scenarios in binary classification with tabular data currently exist. These issues present important topics for future research.
In this work, we used the TableShift benchmark as a proxy, despite the unspecified or unknown shift types for each dataset. Although the benchmark’s authors have made an effort to quantify the amount of shift with regard to each component (covariate, concept, and label shift) for each dataset, they use ad hoc metrics instead of principled ones. In fact, the metrics used for covariate shift and concept shift in their work show almost linear dependency [Gardner et al., 2024], hinting they might be measuring the same phenomenon rather than orthogonal components. The authors correctly point out that measuring the amount of shift is not possible from a finite sample [Gardner et al., 2024]. Indeed, no principled metrics exist for even approximating the shift from a finite sample in any of the components. Quantifying label shift is perhaps the easiest of the three and propositions on how to quantify it have recently been made Sun et al. [2025]. However, shift quantification remains an open research problem.
@@ 6 Conclusion
We presented CBPE, a family of confidence-based performance estimators for model monitoring, which require no access to ground truth labels. Contrary to other confidence-based estimators which can only target model accuracy, CBPE can estimate a binary classifier’s performance with any metric derivable from the confusion matrix, making CBPE more suitable for monitoring classifiers under class imbalance or when some error type is more costly. Also contrary to other estimators, CBPE quantifies uncertainty in the estimates, by yielding a full probability distribution for the estimated metric. We presented theoretical guarantees of its quality under perfect calibration and showed experimentally that its estimation error remains feasible under small calibration error. We hope our work lays the groundwork for future research on unsupervised performance estimation.
Acknowledgements.
@@
Appendix A a Review of Classification Metrics
@@
Actual Positive $TP$ $FN$
Negative $FP$ $TN$
@@ Appendix B Proofs
In the proofs to follow, we use the word "model" to refer exclusively to binary classifiers. We start with the proof of Theorem 1 along with some complementary results. Then, we prove Lemma 4.
@@ B.1 Proof of Theorem 1
For this proof, we will require a slightly altered definition of calibration, where the probability of seeing a positive label is conditioned not only on the confidence score but also on the predicted label. We call this class-conditional calibration with the following formal definition.
@@
Model $f$ is perfectly class-conditionally calibrated within distribution $p_{t}(\textbf{x},y)$ iff.
@@
Since $f$ was assumed to be confidence-consistent, for each $s\in[0,1]$ there exists $a\in\{0,1\}$ such that $P(\hat{Y}=a\mid S=s)=1$. This allows us to simplify
$\sum_{c=0}^{1}P_{t}(Y=1\mid\hat{Y}=c\cap S=s)P(\hat{Y}=c\mid S=s)$ $=s$
@@
where the final equation holds for all pairs $(a,s)$ within the support of $(\hat{Y},S)$. ∎
Let us now consider a model $f$ operating on some target distribution $p_{t}(\boldsymbol{x},y)$. We will use $P_{(x,y)\sim p_{t}(\boldsymbol{x},y)}(\cdot)=P_{t}(\cdot)$, $p_{(x,y)\sim p_{t}(\boldsymbol{x},y)}(\cdot)=p_{t}(\cdot)$, and $\mathbb{E}_{(x,y)\sim p_{t}(\boldsymbol{x},y)}(\cdot)=\mathbb{E}_{t}(\cdot)$ as shorthand notations for any associated probabilities, densities, or expectations respectively. In particular, let $p_{t}(s)$ denote the density function of the distribution of confidence scores induced by model $f$ operating on $p_{t}(\boldsymbol{x},y)$.
@@
Let $(X,Y)$ be an instance drawn from a target distribution $p_{t}(\boldsymbol{x},y)$ and let $\hat{Y}$ be the corresponding prediction made with confidence $S$ by a calibrated and confidence-consistent model $f$. Then,
@@
Consider the probability that the actual label is positive given that it was predicted to be positive with some confidence score $S=s$ and apply the Bayes rule to get
@@
Since our assumptions guarantee the model to be class-conditionally calibrated, we can use Definition 3 and substitute the left side to get
@@
Let $(X,Y)$ be an instance drawn from a target distribution $p_{t}(\boldsymbol{x},y)$ and let $\hat{Y}$ be the corresponding prediction made with confidence $S$ by a calibrated and confidence-consistent model $f$. Then, the following are true
@@
Statement 1. follows directly from Lemma 4 by substituting all instances of the expression $\hat{Y}=1$ with $\hat{Y}=0$ in the proof. Statement 2. also follows from Lemma 4 by applying the complement rule as follows.
@@
Similarly, Statement 3. follows from Statement 1. by applying the complement rule as above. ∎
We are now ready to prove Theorem 1. We split the proof into two parts. In the first part, we show the unbiasedness of the estimator and in the second part, we show the consistency. Again, we sample $n$ points from $p_{t}(\boldsymbol{x},y)$ and collect predictions for them. Assume that there are $n_{+}$ positive and $n_{-}$ negative predictions within the $n$ predictions. Denote the zero vector and a vector of all ones of size $m$ with $\bar{0}_{m}$ and $\bar{1}_{m}$ respectively. Let us start by proving the unbiasedness of our estimators in the following theorem.
@@
Let $f$ be a calibrated and confidence-consistent model operating on some target distribution $p_{t}(\boldsymbol{x},y)$. Then, the following are true
@@
    $\mathbb{E}_{t}\left[\bar{S}^{+}\mid\hat{Y}=\bar{1}_{n_{+}}\right]=\operatorname{TPF}_{t}(\boldsymbol{\mathrm{f}})$
@@
    $\mathbb{E}_{t}\left[\bar{S}^{-}\mid\hat{Y}=\bar{0}_{n_{-}}\right]=\operatorname{FNF}_{t}(\boldsymbol{\mathrm{f}})$
@@
    $\mathbb{E}_{t}\left[1-\bar{S}^{+}\mid\hat{Y}=\bar{1}_{n_{+}}\right]=\operatorname{FPF}_{t}(\boldsymbol{\mathrm{f}})$
@@
    $\mathbb{E}_{t}\left[1-\bar{S}^{-}\mid\hat{Y}=\bar{0}_{n_{-}}\right]=\operatorname{TNF}_{t}(\boldsymbol{\mathrm{f}})$
@@
$\mathbb{E}_{t}\left[\bar{S}^{+}\mid\hat{Y}=\bar{1}_{n_{+}}\right]$ $=\mathbb{E}_{t}\left[\frac{1}{n_{+}}\sum_{i\in\mathrm{I}_{+}}S_{i}\leavevmode\nobreak\ \Bigg{|}\leavevmode\nobreak\ \hat{Y}=\bar{1}_{n_{+}}\right]$
@@
Let $f$ be a calibrated and confidence-consistent model. Then, the following are true
@@
Again, we show only the proof for Statement 1. The proofs for all other statements are completely analogous. Since all $S_{i}$ take values from the interval $[0,1]$, it is trivial that $\mathbb{E}[S_{i}^{2}]\leq\mathbb{E}[S_{i}]$ for all $i$. Since clearly also $\mu:=\mathbb{E}[S_{i}]\in[0,1]$, it is straightforward to see that the expression $\mu(1-\mu)$ is maximized when $\mu=\frac{1}{2}$, yielding a maximum of $\frac{1}{4}$. Using these insights, we can derive an upper bound for the variance of the estimator as follows.
$\operatorname{Var}[\bar{S}^{+}]$ $=\operatorname{Var}\left[\frac{1}{n_{+}}\sum_{i\in I^{+}}S_{i}\right]$
$=\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\operatorname{Var}\left[S_{i}\right]$
$=\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\left(\mathbb{E}[S_{i}^{2}]-\mathbb{E}[S_{i}]^{2}\right)$
$\leq\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\left(\mathbb{E}[S_{i}]-\mathbb{E}[S_{i}]^{2}\right)$
$=\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\mu(1-\mu)$
$\leq\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\frac{1}{4}$
$=\frac{1}{4n_{+}}.$
Since $\underset{n_{+}\rightarrow\infty}{\lim}\frac{1}{4n_{+}}=0$ and $0\leq\operatorname{Var}[\bar{S}^{+}]\leq\frac{1}{4n_{+}}\leavevmode\nobreak\ \forall n_{+}\in\mathbb{Z}^{+}$, it necessarily follows that $\underset{n_{+}\rightarrow\infty}{\lim}\operatorname{Var}[\bar{S}^{+}]=0$. ∎
@@ B.2 Proof of Lemma 2
when $n\to\infty$. Furthermore, the decay rate of the approximation error
$\left|\mathbb{E}\left[\frac{X}{X+Y+an}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[X]+\mathbb{E}[Y]+an}\right|$
is $\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$.
@@
[…]
@@
[…]
@@ Appendix C Additional Algorithms
Data: $X_{metric},\alpha$    (Distribution for a given metric and a confidence level $\alpha$)
Result: $(lower,upper)$  (Bounds for the $1-\alpha$ HDI)
$X_{sorted}\leftarrow sort(X_{metric})$;
$l\leftarrow 0$;
$u\leftarrow|X_{sorted}|-1$;
$tail\_coverage\leftarrow 0$;
$bounds\_not\_found\leftarrow True$;
while $bounds\_not\_found$ do
       $p_{l}\leftarrow X_{sorted}[l][1]$;
       $p_{u}\leftarrow X_{sorted}[u][1]$;
       if $p_{l}<p_{u}$ then
             if $tail\_coverage+p_{l}<\alpha$ then
                   $tail\_coverage\leftarrow tail\_coverage+p_{l}$;
                   $l\leftarrow l+1$;
            else
                   $bounds\_not\_found\leftarrow False$;
             end if
      else
             if $tail\_coverage+p_{u}<\alpha$ then
                   $tail\_coverage\leftarrow tail\_coverage+p_{u}$;
                   $u\leftarrow u-1$
            else
                   $bounds\_not\_found\leftarrow False$
             end if
       end if
end while
$lower\leftarrow X_{sorted}[l][0]$;
$upper\leftarrow X_{sorted}[u][0]$;
return $(lower,upper)$
@@ Appendix D Additional experiments
D.1 Quality of the Confidence Intervals
In this experiment we used the same synthetic data set as in the experiment described in Section 4.1, but with the addition of using the sampled confidence scores to sample a corresponding label for each confidence score from a Bernoulli distribution using the confidence score as a parameter. This way of reverse sampling ensures that the simulated classifier is calibrated by construction although a small calibration error might persist due to sampling effects. After this, we again conducted 10,000 trials for each monitoring window size ranging from 100 to 1,000 and recorded whether the actual value was within a predicted 95% or 90% confidence interval. Finally, we calculated the ratio of these successful trials over all trials, which should be close to 0.95 or 0.9 respectively.
D.2 Simulated covariate shift
As discussed in Section 5, confidence-based estimators are not guaranteed to work under concept shift, which is present in current binary classification benchmarks such as TableShift [Gardner et al., 2024]. Thus, we sought to demonstrate the effectiveness of CBPE in a situation where we could control the precise nature of the shift in the data distribution. For this, we created synthetic data where the concepts remain stable and only covariate shift is present.

### v2

@@ 6 Discussion
It is not immediately obvious how the methods for unsupervised accuracy estimation referenced in the Introduction could be adapted to estimate any other classification metric besides accuracy. As such, CBPE sets the baseline for unsupervised estimation of all binary classification metrics, which can be derived from the confusion matrix. Although CBPE comes with relatively strong theoretical guarantees, it also has its limitations. First and foremost, the theoretical guarantees hinge on the assumption of perfect calibration, which is not achievable in any real-life scenario. However, our experiments show that in most cases, CBPE yields consistent estimates with relatively small estimation errors, as the calibration errors remain small as well. This reduces the problem of performance estimation to a calibration error minimization problem.
Confidence-based performance estimators are known to struggle under concept shift, since even when $p_{s}(\boldsymbol{x},y)$ and $p_{t}(\boldsymbol{x})$ are fully known, the problem of estimating model performance in this setting is fundamentally unidentifiable [garg:2022] and no guarantees for estimation quality can be given. This is because concept shift breaks the stochastic dependency between the features and labels, rendering the confidence scores of a model obsolete in most cases. Also, since calibration error typically increases with covariate shift [ovadia:2019] and label shift [popordanoska:2023], methods for remaining calibrated under these shifts are needed. Unfortunately, no dedicated benchmarks for covariate shift and label shift scenarios in binary classification with tabular data currently exist. These issues present important topics for future research.
In this work, we used the TableShift benchmark as a proxy, despite the unspecified or unknown shift types for each dataset. Although the benchmark’s authors have made an effort to quantify the amount of shift with regard to each component (covariate, concept, and label shift) for each dataset, they use ad hoc metrics instead of principled ones. In fact, the metrics used for covariate shift and concept shift in their work show almost linear dependency [gardner:2024], hinting they might be measuring the same phenomenon rather than orthogonal components. The authors correctly point out that measuring the amount of shift is not possible from a finite sample [gardner:2024]. Indeed, no principled metrics exist for even approximating the shift from a finite sample in any of the components. Quantifying label shift is perhaps the easiest of the three, and propositions on how to quantify it have recently been made [sun:2025]. However, shift quantification remains an open research problem.
@@ 7 Conclusion
We presented CBPE, a family of confidence-based performance estimators for model monitoring, which require no access to ground truth labels. Contrary to other confidence-based estimators, which can only target model accuracy, CBPE can estimate a binary classifier’s performance with any metric derivable from the confusion matrix, making CBPE more suitable for monitoring classifiers under class imbalance or when some error type is more costly. Also, contrary to other estimators, CBPE quantifies uncertainty in the estimates by yielding a full probability distribution for the estimated metric. We presented theoretical guarantees of its quality under perfect calibration and showed experimentally that its estimation error remains feasible under small calibration error. We hope our work lays the groundwork for future research on unsupervised performance estimation.
[…]
Acknowledgements
@@
Appendix A Review of Classification Metrics
@@
Actual Positive TPTP FNFN
Negative FPFP TNTN
@@ Appendix B Proofs
In the proofs to follow, we use the word ”model” to refer exclusively to binary classifiers. We start with the proof of Theorem 1 along with some complementary results. Then, we move on to prove Lemma 2.
@@ B.1 Proof of Theorem 1
For this proof, we require an altered definition of calibration, where the probability of observing a positive label is also conditioned on the predicted label. We call this class-conditional calibration with the following formal definition.
@@
Model $f$ is perfectly class-conditionally calibrated within $p_{t}(\boldsymbol{x},y)$ iff.
@@
where the final step is validated by the fact that $f$ was assumed to be confidence-consistent, and hence for each $s\in[0,1]$ there exists $a\in\{0,1\}$ such that $P(\hat{Y}=a\mid S=s)=1$, which guarantees that the final equation holds for all pairs $(a,s)$ within the support of $(\hat{Y},S)$. ∎
Let us next consider a model $f$ operating on some target distribution $p_{t}(\boldsymbol{x},y)$. Suppose $p_{t}(s)$ denotes the density function of the distribution of confidence scores induced by model $f$ operating on $p_{t}(\boldsymbol{x},y)$.
@@
Let $(X,Y)$ be an instance drawn from a target distribution $p_{t}(\boldsymbol{x},y)$ and let $\hat{Y}$ be the corresponding prediction made with confidence $S$ by a perfectly calibrated and confidence-consistent model $f$. Then,
@@
Consider the probability that the actual label is positive given that it was predicted to be positive with some confidence score $S=s$, and apply the Bayes rule to get
@@
Since our assumptions guarantee the model to be perfectly class-conditionally calibrated, we can use Definition 3 and substitute the left side to get
@@
Let $(X,Y)$ be an instance drawn from a target distribution $p_{t}(\boldsymbol{x},y)$ and let $\hat{Y}$ be the corresponding prediction made with confidence $S$ by a perfectly calibrated and confidence-consistent model $f$. Then, the following are true
@@
Statement 1 follows directly from Lemma 4 by substituting all instances of the expression $\hat{Y}=1$ with $\hat{Y}=0$ in the proof. Statement 2 also follows from Lemma 4 by applying the complement rule as follows.
@@
Similarly, Statement 3 follows from Statement 1 by applying the complement rule as above. ∎
We are now ready to prove Theorem 1. We split the proof into two parts. In the first part, we show the unbiasedness of the estimator, and in the second part, we show the consistency. Again, we sample $n$ points from $p_{t}(\boldsymbol{x},y)$ and collect predictions for them. Assume that there are $n_{+}$ positive and $n_{-}$ negative predictions within the $n$ predictions. Denote the zero vector and a vector of all ones of size $m$ with $\bar{0}_{m}$ and $\bar{1}_{m}$ respectively. Let us start by proving the unbiasedness of our estimators in the following theorem.
@@
Let $f$ be a perfectly calibrated and confidence-consistent model operating on some target distribution $p_{t}(\boldsymbol{x},y)$. Then, the following are true
@@
    $\mathbb{E}_{t}\left[\bar{S}^{+}\mid\hat{Y}=\bar{1}_{n_{+}}\right]=\operatorname{TPF}_{t}(f)$
@@
    $\mathbb{E}_{t}\left[\bar{S}^{-}\mid\hat{Y}=\bar{0}_{n_{-}}\right]=\operatorname{FNF}_{t}(f)$
@@
    $\mathbb{E}_{t}\left[1-\bar{S}^{+}\mid\hat{Y}=\bar{1}_{n_{+}}\right]=\operatorname{FPF}_{t}(f)$
@@
    $\mathbb{E}_{t}\left[1-\bar{S}^{-}\mid\hat{Y}=\bar{0}_{n_{-}}\right]=\operatorname{TNF}_{t}(f)$
@@
$\mathbb{E}_{t}\left[\bar{S}^{+}\mid\hat{Y}=\bar{1}_{n_{+}}\right]$ $=\mathbb{E}_{t}\left[\frac{1}{n_{+}}\sum_{i\in\mathrm{I}_{+}}S_{i}~\Bigg|~\hat{Y}=\bar{1}_{n_{+}}\right]$
@@
Let $f$ be a perfectly calibrated and confidence-consistent model. Then, the following are true
@@
Again, we show only the proof for Statement 1. The proofs for all other statements are completely analogous. The variance of $\bar{S}^{+}$ is
$\operatorname{Var}[\bar{S}^{+}]=\operatorname{Var}\left[\frac{1}{n_{+}}\sum_{i\in I^{+}}S_{i}\right]=\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\operatorname{Var}\left[S_{i}\right]=\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\left(\mathbb{E}[S_{i}^{2}]-\mathbb{E}[S_{i}]^{2}\right).$
Since all $S_{i}$ take values from the interval $[0,1]$, it is trivial that $\mathbb{E}[S_{i}^{2}]\leq\mathbb{E}[S_{i}]$ for all $i$. Since clearly also $\mu:=\mathbb{E}[S_{i}]\in[0,1]$, it is straightforward to see that the expression $\mu(1-\mu)$ is maximized when $\mu=\frac{1}{2}$, yielding a maximum of $\frac{1}{4}$. Using these insights, we can derive an upper bound for the variance of the estimator as follows.
$\operatorname{Var}[\bar{S}^{+}]\leq\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\left(\mathbb{E}[S_{i}]-\mathbb{E}[S_{i}]^{2}\right)=\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\mu(1-\mu)\leq\frac{1}{n_{+}^{2}}\sum_{i\in I^{+}}\frac{1}{4}=\frac{1}{4n_{+}}.$
Since $\underset{n_{+}\rightarrow\infty}{\lim}\frac{1}{4n_{+}}=0$ and $0\leq\operatorname{Var}[\bar{S}^{+}]\leq\frac{1}{4n_{+}}$, it follows that $\underset{n_{+}\rightarrow\infty}{\lim}\operatorname{Var}[\bar{S}^{+}]=0$. ∎
@@ B.2 Proof of Lemma 2
when $n\to\infty$. Furthermore, the decay rate of the approximation error is $\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$.
@@
$\frac{X-\mathbb{E}[X]}{\sqrt{\operatorname{Var}(X)}}$ $=\frac{X-\mathbb{E}[X]}{\sqrt{n_{1}\sigma_{\bar{X}_{i}}^{2}}}=\frac{1}{\sqrt{n_{1}}}\cdot\frac{X-\mathbb{E}[X]}{\sigma_{\bar{X}_{i}}}\xrightarrow{~D~}\mathcal{N}(0,1),$
$\frac{Y-\mathbb{E}[Y]}{\sqrt{\operatorname{Var}(Y)}}$ $=\frac{Y-\mathbb{E}[Y]}{\sqrt{n_{2}\sigma_{\bar{Y}_{j}}^{2}}}=\frac{1}{\sqrt{n_{2}}}\cdot\frac{Y-\mathbb{E}[Y]}{\sigma_{\bar{Y}_{j}}}\xrightarrow{~D~}\mathcal{N}(0,1),$
@@
$\frac{X}{X+Y+an}\xrightarrow{~p~}\frac{\mathbb{E}[X]}{\mathbb{E}[X]+\mathbb{E}[Y]+an},$
@@ Appendix C Additional Algorithms
[…]
2:$(lower,upper)$  (Bounds for the $1-\alpha$ HDI)
[…]
8:while bounds_not_foundbounds\_not\_found do
9:  $p_{l}\leftarrow X_{sorted}[l][1]$
10:  $p_{u}\leftarrow X_{sorted}[u][1]$
11:  if $p_{l}<p_{u}$ then
12:   if $tail\_coverage+p_{l}<\alpha$ then
13:     $tail\_coverage\leftarrow tail\_coverage+p_{l}$
14:     $l\leftarrow l+1$
15:   else
16:     $bounds\_not\_found\leftarrow False$
17:   end if
18:  else
19:   if $tail\_coverage+p_{u}<\alpha$ then
20:     $tail\_coverage\leftarrow tail\_coverage+p_{u}$
21:     $u\leftarrow u-1$
22:   else
23:     $bounds\_not\_found\leftarrow False$
24:   end if
25:  end if
26:end while
[…]
@@ Appendix D Additional experiments
D.1 Simulated covariate shift
As discussed in Section 6, confidence-based estimators are not guaranteed to work under concept shift, which is present in current binary classification benchmarks such as TableShift [gardner:2024]. Thus, we sought to demonstrate the effectiveness of CBPE in a situation where we could control the precise nature of the shift in the data distribution. For this, we created synthetic data where the concepts remain stable and only covariate shift is present.